A Michelson interferometer is irradiated with light of wavelength 633 nm (He-Ne laser). When one of the mirrors is moved, 800 fringe pairs (each fringe pair is an adjacent dark and bright line) pass by a fixed point in the viewing pattern. How far was the mirrow moved?
Answer #1

The mirror was moved

d = m*(wavelength)/2

where

wavelength = 633 nm
m = 800

Thus, the mirror was mnoved by

d = 800*633/2 nm

d = 253200 nm or 0.253 mm [ANSWER]

DONE! It was nice working with you!
Answer #2

in Michelson interferometer,

the moved by the mirror is ,

(x2-x1)=N*lambda/2

=800*633*10^-9/2

=2.532*10^-4 m

=0.2532 mm ..............is answer